A DVD costs £12.95. In a sale all prices are reduced by 3/10 what is the sale price of the DVD
1
Expert's answer
2020-01-20T10:18:50-0500

If price is reduced by 3/10 (and the whole price is 10/10=1), the sale price is 10/10 - 3/10 = 7/10.

Let x be the sale price of the DVD

Thus, using a proportion we may see

12.95 - 1

x - 0.7

Thus, x = (12.95*0.7)/1 = 9.065

That means the sale price of the DVD is £9.065 (or £9.07).